Hedging Decision. Indiana Company expects to receive 5 million euros in one year from exports.

It can use any one of the following strategies to deal with the exchange rate risk. Estimate the dollar cash flows received as a result of using the following strategies:

a) unhedged strategy

b)money market hedge

c)option hedge

The spot rate of the euro as of today is $1.10. Interest rate parity exists. Indiana Company uses the forward rate as a predictor of the future spot rate. The annual interest rate in the U.S. is 8% versus an annual interest rate of 5% in the euro zone. Put options on euros are available with an exercise price of $1.11, an expiration date of one year from today, and a premium of $.06 per unit. Estimate the dollar cash flows it will receive as a result of using each strategy. Which hedge is optimal?
Answer #1

Future Spot Rate based on Interest Rate Parity = Spot Rate × (Interest Rate in US/Interest Rate in Euro)

= $1.1 × [(1+0.8)/(1+0.5)] = $1.1314

Dollar Cash Flow under different strategies:

(a) Unhedged Strategy

= €5,000,000 ×1.1314 = $5,657,000

(b) Money Market Hedge

Steps:

Now,

(1) Borrow 5,000,000/1.05 = €4,761,904.76 for 1 year at 5%

(2) Convert €4,761,904.76 into $ at spot rate of $1.1 and receive $5,238,095.236

(3) Invest $5,238,095.236 for 1 year at 8%

1 year later

(4) Receive €5,000,000 from exports

(5) Repay borrowed loan of €4,761,904.76 along with 5% interest i.e. 4,761,904,76×1.05 = €5,000,000

(6) Recieve proceeds from investments i.e. 5,238,095.236×1.08 = $5,657,142.85

Therefore, Net Dollar Cash Flow = $5,657,142.85

(c) Options Hedge

Buy Put options for €5,000,000 of strike price $1.11 at a premium of $0.06

Therefore, Cash Outflow TODAY = 5,000,000×0.06 = $300,000

As Interest Rate Parity exists, future spot rate will be $1.1314, whereas strike price of put option is $1.11. Therefore, it is better NOT TO EXERCISE THE OPTION.

Therefore, Net Dollar Cash Flow = (5,000,000×$1.1314) - PREMIUM PAID

PREMIUM IS PAID TODAY AND CASH FLOWS ARE RECEIVED AFTER 1 YEAR. THEREFORE, 1 YEAR INTEREST SHOULD BE ADDED TO PREMIUM PAID.

= $5,657,000 - [$300,000×1.08]

= $5,333,000

Net Dollar Cash Flow will be $5,333,000.

Net Dollar Cash Flow is Highest in Money Market Hedge option. So, it is the optimal option.